OpenStudy (anonymous):

2. Simplify the given expression to rational exponent form, justify each step by identifying the properties of rational exponents used.

OpenStudy (anonymous):

I dont want the answer, I just want to know how to start this off.

OpenStudy (anonymous):

Start with \[ \sqrt[n]{x} = x^{1/n} \]

OpenStudy (anonymous):

So \[ \sqrt[3]{x^{-6}} = \left(x^{-6}\right)^{1/3} \]

OpenStudy (anonymous):

Does that make sense?

OpenStudy (anonymous):

Well, how did you get rid of the radical?

OpenStudy (anonymous):

This is the pattern \[ \sqrt[\color{blue}{n}]{\color{red}{x}} = \color{red}{x}^{1/\color{blue}{n}} \]

OpenStudy (anonymous):

So \[ \sqrt[\color{blue}{3}]{\color{red}{x^{-6}}} = \left(\color{red}{x^{-6}}\right)^{1/\color{blue}{3}} \]

OpenStudy (anonymous):

Can you understand the pattern?

OpenStudy (anonymous):

I think so. Thanks.

OpenStudy (anonymous):

What about that numerator 1 over the radical equation part.

OpenStudy (anonymous):

Well \[ \frac{1}{\color{red}x} = \color{red}x^{-1} \]

OpenStudy (anonymous):

But first we need to simplify \[ ({x^a})^b = x^{a\times b} \]

OpenStudy (anonymous):

So \[ \left(x^{-6}\right)^{1/3} = x^{-6\times 1/3} \]

OpenStudy (anonymous):

And \(-6\times 1/3 = -2\) so \[ x^{-6\times 1/3} = x^{-2} \]

OpenStudy (anonymous):

We have simplified it a bit: \[ \frac{1}{\sqrt[3]{x^{-6}}} =\frac{1}{x^{-2}} \]

OpenStudy (anonymous):

Does this part make sense?

OpenStudy (anonymous):

I see. Yes, it does. I just need to learn the pattern.

OpenStudy (anonymous):

Can you do the rest?

OpenStudy (anonymous):

Yes. Thank you ^.^

OpenStudy (anonymous):

The answer would be x^2 right?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Thank you for your help.